OpenStudy (hey.itsbri):

create an equation from these roots: 2√3, -2√3,1,-3

OpenStudy (mhchen):

Do you know what the root of an equation is?

likeabossssssss (likeabossssssss):

welcome to os and plz close this :)

OpenStudy (naka354):

(x-1)*(x+3)*(x-2√3)*(x+2√3) = 0 (x^2 + 2x - 3)*(x^2 - 12) = 0 x^4 + 2x^3 - 15x^2 - 24x + 36 = 0